OpenStudy (baldymcgee6):

Regarding differentiating... Is it always possible to rearrange a quotient into a product and apply the product rule?

OpenStudy (anonymous):

usually you can be simply doing the quotient rule is much easier in most cases

OpenStudy (anonymous):

*but

OpenStudy (anonymous):

ya, its always possible, but if you do so, sometimes it becomes too lenthy ;)

OpenStudy (anonymous):

if it comes to differentiate 1/x^2 you can write it as, (x)^-2

OpenStudy (baldymcgee6):

i'm aware of how to re-arrange functions, I was just curious to know if there is ever a case where we cannot use the product rule on a quotient

OpenStudy (anonymous):

no because you prove it via the product rule it is always the case that \(\frac{f}{g}=f\times \frac{1}{g}\) but why you would want to work with \(\frac{1}{g}\) is not always clear